Organic Chemistry

What s basic difference b/w aromatic , antiaromatic and nonaromatic

anoop  ranga
14 Years agoGrade

1 Answer

Rahul Kumar
14 Years ago

Hi Anoop

aromatic compound follows (4n+2)pi rule.where n=0,1,2,3,...........etc. ex. benzene has 6 pi electron. that is (4*1+2)=6. it obey (4n+2) pi rule. hence it is aromatic.

anti aromatic follows 4n pi rule. where n=0,1,2,3,..........etc. ex. butadine has 4 electron. that is it obey 4*1=4 rule.

the compound which do not follow above two rules they are non aromatic.